A recent study reports that 2.5% of elementary children are allergic to peanuts. A random sample of 200 elementary school children is collected to test this.

Compute the probability that proportion of children in the sample of 200 who are allergic to peanuts is 3% or less. ( < .03).

Answer by stanbon(75887)

A recent study reports that 2.5% of elementary children are allergic to peanuts. A random sample of 200 elementary school children is collected to test this.
Compute the probability that proportion of children in the sample of 200 who are allergic to peanuts is 3% or less. ( < .03).
----------------------------------
z(0.03) = (0.03-0.025)/sqrt[0.025*0.975/200] = 0.4529
----
P(p-hat < 0.03) = P(z < 0.4529) = normalcdf(-100,0.4529) = 0.6750
========================
Cheers,
Stan H.
============